OpenStudy (anonymous):

The first part of the FTC states that if F(x) = integral of f(t) from a to x, then F'(x) = f(x). Using these facts and the properties of the definite integral find the derivative of ....ill write the integral using the equation tool

OpenStudy (anonymous):

\[h(x) = \int\limits_{\sin^2x}^{x^2} tsin \sqrt{t} dt\]

OpenStudy (anonymous):

Its hard to see but the bounds say: lower bound = sin^2 x and the upper bound is x^2

OpenStudy (abb0t):

It is the derivative times the function according to FToC.

OpenStudy (abb0t):

both the upper and lower boundry derivatives times f(g(x)) for both. Since you have two functions, you are basically doing it twice.

OpenStudy (abb0t):

where g(x) is your boundries of integration functions.

OpenStudy (anonymous):

So i separted the functions to \[\int\limits_{0}^{x^2}tsin \sqrt{t}dt - \int\limits_{0}^{si^2x}t \sin \sqrt{t}dt\]

OpenStudy (anonymous):

So h(x) = f(g(x)) and \[f'(x) = xsin \sqrt{x}\]

OpenStudy (abb0t):

If I remember correctly: \[\int\limits_{0}^{v(x)}f(t)dt+\int\limits_{0}^{u(x)}f(t)dt\]

OpenStudy (anonymous):

Ok so i did that part correctly. What should i do next

OpenStudy (abb0t):

\[\frac{ d }{ dx }=-v'(x)f(v(x))+u'(x)f(u(x))\]

OpenStudy (abb0t):

for the first integral, the v(x) should be below the zero. \[\int\limits_{v(x)}^{0}\]

OpenStudy (anonymous):

Oh so the lower bound should be x^2 and the upper bound is 0.

OpenStudy (abb0t):

Yeah. That's why you get a -v'(x)

OpenStudy (anonymous):

ok and now its \[-x^2(f(x^2)) + \sin^2x(f(\sin^2x))\]

OpenStudy (abb0t):

Yeah.

OpenStudy (abb0t):

well, that's not actually the derivative. so not exactly.

OpenStudy (abb0t):

\[v'(x) \] and \[u'(x)\]

OpenStudy (anonymous):

ok so since v(x) = x^2 v'(x) = 2x. u(x) = sin^2x so u'(x) = 2sinx(cosx)

OpenStudy (abb0t):

yep. but the original function is what gets plugged in as t.

OpenStudy (abb0t):

you just multiply by the derivative.

OpenStudy (anonymous):

so its -2x(f(x^2)) + 2sinxcosx ( f(sin^2 x))

OpenStudy (abb0t):

yes. correct.

OpenStudy (anonymous):

so we plug the x^2 and sin^2 x into the tsin*radical t equation??

OpenStudy (abb0t):

yuup.

OpenStudy (anonymous):

\[f(x) = x^2\sin \sqrt{x^2}\] \[f(\sin^2x) = \sin^2x* \sin \sqrt{x^2} = sinx^3\sqrt{x^2}\]

OpenStudy (anonymous):

for the 1st equation i mean f(x^2) not f(x)

OpenStudy (abb0t):

You can't just combine both sin's because they are not the same. One is sin(x) and the other is sin(x^1/2)

OpenStudy (abb0t):

and you also plug in sin^2(x) under the square root.

OpenStudy (anonymous):

oh right its \[\sin^2x \sqrt{\sin^2x} = \sin^2x * sinx = \sin^3 x\]

OpenStudy (anonymous):

so i cant combine the exponents for this one?

OpenStudy (abb0t):

No. Both sin's are different. However, if you had \[\sin(2x)\sin(2x) = \sin^2(2x)\]

OpenStudy (anonymous):

I hope this is correct now

OpenStudy (anonymous):

\[-2x(x^2\sin \sqrt{x^2}) + 2sinxcosx( \sin^2x \sqrt{\sin^2x})\]

OpenStudy (anonymous):

i should distribute this now right?

OpenStudy (anonymous):

nothing cancels out so i guess i should simplify this as much as possible

OpenStudy (abb0t):

yes. u should to make it look neater. but essentially, thats ur answer.

OpenStudy (anonymous):

when u distribute the -2x to the first parenthesis do u get \[-2x^3(-2xsin \sqrt{x^2})\]

OpenStudy (anonymous):

that doesnt look right. What should it be

OpenStudy (abb0t):

\[-2x^2(x^2\sin(\sqrt{x^2}+2\sin(x)\cos(x)\sin(\sqrt \sin^2(x))\]

OpenStudy (abb0t):

the first term should be 2x. not 2x^2

OpenStudy (anonymous):

why do u distribute the -2x to everything. I thought it was only for the 1st parenthesis

OpenStudy (abb0t):

I didn't distribute the 2x to everything. I'm missing a parenthesis. here: \[-2x(x^2\sin(\sqrt {x^2}) )+2\sin(x)\cos(x)(\sin^2(x)(\sqrt{\sin^2(x}))\]

OpenStudy (anonymous):

ok yes. Thats also what i got. Its hard to simplify that :/

OpenStudy (abb0t):

and you know by trig identities, \[2\sin(x)\cos(x) = \sin(2x)\] if you wish to take it further. but I don't think you need to.

OpenStudy (anonymous):

ok so u think thats enough simlification?

OpenStudy (abb0t):

wait, nevermind. i think that trig identity is wrong. haha.

OpenStudy (abb0t):

you might want to double check my trig identity.

OpenStudy (anonymous):

i googled it now. U are correct

OpenStudy (abb0t):

i am sure that for Calculus, that asnwer is good enough. simplifying further is algebra, and you are past that. it's assumed that you know your algebra tricks very well by now.

OpenStudy (anonymous):

OK! THANK YOU 100 TIMES OVER! YOU are very patient to explain everything to me! Thanks bro. I really needed the help!

OpenStudy (anonymous):

I understand teh calculus. Lord help me in algebra lol!

OpenStudy (abb0t):

Calculus is just advance algebra. b/c as you may have noticed, it's the algebra that stoops most people. The calculus is relatively easy. You have formulas that u use to manipulate algebraically.

OpenStudy (anonymous):

Yeah. The easy part is definitly the calculus part b/c its just understanding the concept. The algebra will just need some brushing up on

OpenStudy (anonymous):

Thanks! This was more than enough help. Ur great man.

OpenStudy (abb0t):

fa sho. best of luck, mate. Make sure you get integrals well cuz it gets way tougher in Calc-II.

OpenStudy (anonymous):

Ok thanks for the tip! I'll make sure I get it down! Thanks. Peace.